Circular No. 7874

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1908+094 M. P. Rupen, V. Dhawan and A. J. Mioduszewski, National Radio

Astronomy Observatory, report: “We have detected a possible radio counterpart to the black-hole candidate XTE J1908+094. Observations with the Very Large Array on Mar. 21 and 22 show a radio source (flux 1.1 ± 0.1 mJy at 4860 MHz, 0.84 ± 0.08 mJy at 8460 MHz) at α = 19^{h}08^{m}53^{s}.077, δ = +9^{o}23′04′′.90 (J2000.0, uncertainty ∼ 0′′.1 in each coordinate), within the 17′′ error box reported from BeppoSAX (IAUC 7873) and ∼ 8′′ from its center. The radio spectrum is consistent with this identification, although no variability has yet been seen, and there is no obvious extension in the images (resolution 0′′.20). Further monitoring is underway.”

XTE J1751–305 Z. Wang, D. Chakrabarty and J. L. Elliot, Massachusetts Institute

of Technology, report: “We have identified a possible infrared counterpart for the new accreting millisecond pulsar XTE J1751–305 (IAUC 7867, 7870, 7872). We obtained infrared observations of the pulsar field with the Magellan Consortium’s 6.5-m Baade telescope at the Las Campanas Observatory on Apr. 9. Only one star lies within the provisional 0′′.5 error circle measured by the Chandra X-Ray Observatory on Apr. 10 (C. B. Markwardt, private communication), down to a limiting magnitude of J = 22. The magnitude of this star was J = 19.2 ± 0.2. However, based on the observed density of stars in this Galactic center field, there is a significant chance coincidence probability of 20 percent. The preliminary position for the infrared candidate is α = 17^{h}51^{m}13^{s}.51, δ = −30^{o}37′23′′.1 (equinox 2000.0, 0′′.2 uncertainty). Further observations are required to verify this candidate by searching for 42-min orbital variability and fading as the x-ray source declines.”

SUPERNOVA 2002bz IN MCG +05-34-33 T. Matheson, S. Jha, P. Challis and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002bz (cf. IAUC 7866), obtained by P. Berlind on Apr. 5.36 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type Ia supernova near maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 11 090 km/s for the host galaxy, the supernova expansion velocity is ∼ 10 900 km/s for Si II (rest 635.5 nm) and 14 000 km/s for Ca II (rest 395.1 nm). The spectral-feature supernova age (Riess et al. 1997, A.J. 114, 722) is 3 ± 2 days before maximum light.

2002 April 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Brian G. Marsden